OpenStudy (anonymous):

sinx(tan^2x)(1-sin^2x)

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

simplify the expression

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

\[\Large sinx(\tan^2x)(1-\sin^2x)\] Our original equation Do you know any trig identities?

OpenStudy (anonymous):

Yeah but I'm mixing them up :/

OpenStudy (anonymous):

Ok it's ok, we'll start with a simple one that relates to this equation, you're familar with \[\Large \cos ^{2}(x)+\sin ^{2}(x)=1\] yes?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

Well get the cos^2(x) by itself in that equation I just provided

OpenStudy (anonymous):

1-sin^2(x)

OpenStudy (anonymous):

Ok and so now you can replace that into your equation

OpenStudy (anonymous):

Next question, can you rewrite tan in terms of sin and cos, in other words what is the tan really using the sin and cos?

OpenStudy (anonymous):

isn't that with just tanx=(sinx/cosx)? or does tan^2(x)= ((sin^2(x))/(cos^2(x))?

OpenStudy (loser66):

yes

OpenStudy (anonymous):

Yes you are correct @love4bizarre , our jo here is to guide someone to the solution rather than just give it up @Loser66 , it's just a heads up I've seen people get banned like that

OpenStudy (anonymous):

\[\Huge \tan ^{2}(x)=\frac{\sin^{2}(x)}{\cos^{2}(x)}\]

OpenStudy (anonymous):

So basically the new equation is \[\Large sinx \times(\frac{ \sin^2(x) }{ \cos^2(x) })\times(\cos^{2}(x)) \]

OpenStudy (anonymous):

Now you just simplify what you can and multiply the rest. Tell me what you get.

OpenStudy (anonymous):

I got sin^3(x)

OpenStudy (anonymous):

is that okay?

OpenStudy (anonymous):

Yep!, you got it!

OpenStudy (anonymous):

Thank you for explaining it so well :)